# Admissible covers and stable maps

### v1

@@ 7.8. Further simplification of the wall-crossing formula for point insertions
(7.4$)=\sum_{\begin{subarray}{c}b,\Gamma,\underline{\mu}\end{subarray}}\mathrm{Hur}^{X}(\underline{\mu})\cdot\prod_{i=1}^{n}{\mathfrak{z}}(\mu^{i})\mathrm{Hur}^{{\mathbb{P}}^{1}}\left(\mu^{i},\eta^{i},\frac{(-(2))^{b_{i}}}{b_{i}!}\right)\cdot I_{g_{i},\eta^{i}}(\tau_{k_{i}})$
@@ 8.1. Operator expressions
for a character of the symmetric group $\chi^{\lambda}_{\eta}$ and the dimension of the representation $\dim\lambda$ associated to a partition $\lambda$. For positive integers $\eta_{i}$, recall the operator expression for Hodge integrals,
@@ A.1. Implementation and methodology
The computational verification was implemented using the admcycles [DSvZ21], with code available at

### v2

@@ 7.8. Further simplification of the wall-crossing formula for point insertions
(7.4$)=\sum_{\begin{subarray}{c}b,\Gamma,\underline{\mu}\end{subarray}}\mathrm{Hur}^{X}(\underline{\mu})\cdot\prod_{i=1}^{n}{\mathfrak{z}}(\mu^{i})\mathrm{Hur}^{{\mathbb{P}}^{1}}\left(\mu^{i},\eta^{i},\frac{(-(2))^{b_{i}}}{b_{i}!}\right)\cdot I_{g_{i},\eta^{i}}(\tau_{k_{i}}).$
@@ 8.1. Operator expressions
for a character of the symmetric group $\chi^{\lambda}_{\eta}$ of an element in the conjugacy class $\eta$ acting on the representation associated to $\lambda$, and $\dim\lambda$ is the dimension of the representation. For positive integers $\eta_{i}$, recall the operator expression for Hodge integrals,
@@ A.1. Implementation and methodology
The computational verification was implemented using the admcycles package [DSvZ21], with code available at